### v1

# Energy shifts and broadening of excitonic resonances in electrostatically-doped semiconductors

Abstract
Tuning the density of resident electrons or holes in semiconductors through electrostatic doping provides crucial insight into the composition of the many excitonic complexes that are routinely observed as absorption or photoluminescence resonances in optical studies. Moreover, we can change the way these resonances shift and broaden in energy by controlling the quantum numbers (e.g., spin and valley) of the resident carriers with applied magnetic fields and doping levels, and by selecting the quantum numbers of the photoexcited or recombining electron-hole (e-h) pair through optical polarization. Here, we discuss the roles of distinguishability and optimality of excitonic complexes, showing them to be key ingredients that determine the energy shifts and broadening of optical resonances in charge-tunable semiconductors. A distinguishable e-h pair means that the electron and hole undergoing photoexcitation or recombination have quantum numbers that are not shared by any of the resident carriers. An optimal excitonic complex refers to a complex whose particles come with all available quantum numbers of the resident carriers. Based on the carrier density, magnetic field, and light polarization, all optical resonances may be classified as either distinct or indistinct depending on the distinguishability of the e-h pair, and the underlying excitonic complex can be classified as either optimal or suboptimal. The universality of these classifications, inherited from the fundamental Pauli exclusion principle, allows us to understand how optical resonances shift in energy and whether they should broaden as doping is increased. This understanding is supported by conclusive evidence that the broadening and decay of optical resonances cannot be simply attributed to enhanced screening when resident carriers are added to a semiconductor. Finally, applying the classification scheme in either monolayer or moiré heterobilayer semiconductor systems, we relate the energy shift and amplitude of the neutral exciton resonance to the compressibility of the resident carrier gas.
I Introduction
Optical absorption and emission processes in electrostatically-doped semiconductors tell us a great deal about the pervading role of the Pauli exclusion principle. One ramification of the Pauli principle in semiconductors is the quantum-mechanical distinguishability of particles in excitonic complexes in terms of their valley and spin quantum numbers [1, 2, 3, 4, 5, 6, 7, 8]. In addition, the Pauli principle restricts the permissible types of recombination channels of dark excitons and trions [9, 10, 11, 12], the way hot excitons relax in energy [13, 14, 15, 16, 17, 18], and how impurities assist recombination processes [19]. However, when it comes to the relation between electrostatic doping and the broadening of excitonic optical resonances, or between electrostatic doping and the energy shifts of these resonances [20, 21, 22, 23, 24, 25, 26, 27], the Pauli exclusion principle is often given a lesser role and other mechanisms are invoked to explain such physics. To date, no proposed mechanism can satisfactorily or self-consistently explain the broadening and energy shifts of all optical resonances in charge-tunable semiconductors, including theoretical frameworks based on exciton-polarons [28, 29, 30, 31, 32], correlated trions (i.e. “tetrons”) [35, 33, 34, 36, 37, 38, 39], intervalley plasmons [40, 41, 42, 40], or arguments that rely on band filling and screening [43].
To explain all experimental observations self-consistently and to inform on the essential role played by the Pauli exclusion principle, we propose in this work that all excitonic optical resonances should be classified as being either distinct or indistinct, and that their underlying excitonic complexes should be classified as being either optimal or suboptimal. We will show that this classification scheme provides a useful framework to understand whether and how various excitonic resonances shift and broaden in energy as the charge density is varied.
Distinct versus indistinct optical resonances: When a semiconductor is doped with resident electrons in the conduction band (CB) or resident holes in the valence band (VB), there is an important difference between optical transitions that involve distinguishable versus indistinguishable e-h pairs. A distinguishable pair means that the electron and hole have unique quantum numbers that are not shared by any of the resident carriers. For example, when the electron and hole are photoexcited into unoccupied valleys [44]. Distinct (indistinct) optical resonances arise from optical transitions of distinguishable (indistinguishable) e-h pairs. To streamline the discussion in the rest of this work, we will also refer to resident carriers as indistinguishable if they share quantum numbers with the electron or hole in the pair, and as distinguishable otherwise.
Optimal versus suboptimal excitonic complexes: We say that an excitonic complex is optimal if there are no resident carriers in the system with which the complex could bind to form a different lower-energy complex with more particles. Since particles with similar quantum numbers cannot be part of the same bound complex [45], the electrons (or holes) of an optimal excitonic complex are mutually distinguishable and they include all possible valley-spin species of resident electrons (or holes) at the edge of the CB (VB). This definition implies that the ground-state excitonic complex is always optimal. However, we will show that other complexes can be optimal as well.
We will employ these definitions and present an overarching framework by which one can directly analyze the measured behavior of all optical resonances in charge-tunable semiconductors. Focusing on transition-metal dichalcogenide (TMD) monolayers and moiré heterobilayers by using the experimental findings of Refs. [7, 8, 11, 20, 21, 22, 23, 24, 25, 26, 27], we identify five important understandings and universal observations.
  1. 1.
    Distinct resonances of optimal excitonic complexes neither broaden nor decay when resident carriers are added to the semiconductor. Examples include certain types of trions in valley-polarized monolayers, 6-particle hexcitons, or intralayer excitons in moiré heterobilayers. These resonances only show a steady energy redshift that is likely caused by a small difference between renormalizations of the band gap and binding energies. There is no broadening or decay in this case because the absorption or recombination processes do not conflict with Pauli exclusion and therefore there is no need to shake-up the Fermi sea (i.e., to rearrange the resident carriers).
  2. 2.
    Indistinct resonances of optimal excitonic complexes do broaden and slowly shift in energy due to shakeup processes when the charge density increases. Examples include certain trions in zero magnetic field.
  3. 3.
    While Coulomb screening by resident carriers gives rise to the small difference between renormalizations of the band gap and binding energies, we present conclusive evidence that this screening does not lead to the decay and broadening of any resonances in the optical spectra.
  4. 4.
    With increasing charge density, the energy blueshift and broadening of the neutral-exciton absorption resonance are especially strong in the presence of distinguishable resident carriers. Their shakeup leads to a stronger blueshift than the blueshift caused by Pauli blocking (band filling). That is, stronger in comparison to the increase of the optical energy gap when indistinguishable resident carriers occupy low energy states in the photoexcited valley.
  5. 5.
    The energy shift and amplitude of the neutral exciton resonance are related to the compressibility of the resident carrier gas. We explain the unique behavior of this resonance when the resident carriers form an incompressible many-body state, including integer filling of Landau levels in a monolayer subjected to a strong magnetic field or fractional filling of a moiré lattice at certain doping levels.
Before embarking, it is important to state that our analysis is valid in cases that the neutral exciton binding energy is much larger than the Fermi energy or cyclotron energy when applying a magnetic field, $\varepsilon_{X}\gg E_{F},\,\hbar\omega_{c}$, meaning that the effective Bohr radius of the exciton is smaller than the cyclotron radius or the average distance between resident carriers.
Organization of the paper
Section II outlines the classification of optical resonances and excitonic complexes in MoSe2 and WSe2 monolayers based on their distinctiveness and optimality. Section III discusses various experimental results and then focuses on the universal energy redshift of distinct resonances of optimal excitonic complexes when the charge density increases in monolayers and moiré heterobilayers. We then show that screening by resident carriers is not leading to the decay of excitonic complexes and that band filling is not the primary source of energy blueshift. Section IV deals with shakeup processes, first focusing on the broadening and energy shifts of indistinct resonances of optimal excitonic complexes, and then focusing on the energy blueshift of photoexcited neutral excitons under various conditions. Section V deals with the relation between the compressibility of the resident carrier gas and the behavior of neutral exciton resonances in monolayers and moiré heterobilayers. A summary is given in Sec. VI.
II classification of optical resonances and excitonic complexes
Figure 1 shows examples of exciton and trion optical transitions in electrostatically-doped MoSe2 monolayers, and how these transitions and the underlying excitonic complexes are classified. We choose to start with MoSe2 for its simplicity, where optical transitions are between the upper VBs and lower CBs in both the $K$ and $-K$ valleys. The upper spin-orbit-split CBs (not shown) are assumed to be unpopulated and hence play no role. The carrier populations in the upper VBs and lower CBs are controlled by doping level and by applied magnetic fields $B$. Without loss of generality, we discuss the resonances and exciton complexes as arising from optical transitions in the $K$ valley (i.e., with $\sigma_{+}$ circularly polarized light).
Figures 1(a) and (b) show the photoexcitation in electron-rich conditions. Both exciton and trion resonances are said to be indistinct because the photoexcited electron is promoted to a CB that is already occupied by resident electrons having the same quantum numbers (spin and valley). The neutral exciton in Fig. 1(a) is suboptimal because there are opposite-spin electrons in the $-K$ CB with which this exciton could bind to form a lower-energy complex with more particles (i.e., a trion). Conversely, the negative trion shown in Fig. 1(b) is optimal because the photoexcited pair binds to an electron from the valley at $-K$, and there are no other kinds of resident electrons with which the pair can bind. The trion is depicted as the binding between the photoexcited e-h pair and particle-hole excitation of the Fermi sea in the $-K$ valley, and the entire complex is known as a tetron, wherein the trion moves together and is correlated with the CB hole in the Fermi sea around the trion [35, 33, 34, 36]. Figures 1(c) and (d) show the respective physics in a hole-rich MoSe2 monolayer.
Figure 1: Photoexcitation of the $K$ valley in an electrostatically-doped MoSe2 monolayer. Valleys with solid (dashed) lines host spin-up (spin-down) electrons. The diagrams describe neutral excitons ($X^{0}$) and trions ($X^{\pm}$). For the latter, the photoexcited e-h pair binds to a charge particle from the $-K$ valley, shown by a particle-hole excitation of the Fermi sea. The diagrams in (e)-(f) describe photoexcitation under a strong magnetic field $B$, such that Landau quantization of VB holes (CB electrons) happens solely at the $K$ ($-K$) valley. All optical transitions except for (f) are indistinct resonances, because the photoexcited electron or hole is injected into already-occupied bands.
Applying a magnetic field $B$ can change the classifications if $B$ is strong enough to fully valley-polarize the Fermi sea of resident carriers. In the hole-rich case, shown in Fig. 1(e), the absence of resident VB holes in the $-K$ valley (with which the neutral exciton could otherwise bind) renders $X^{0}$ an optimal complex, although its resonance remains indistinct. In the electron-rich case, shown in Fig. 1(f), the absence of resident CB electrons at the $K$ valley of the photoexcited electron causes the resonance of this optimal trion to become distinct.
The excitonic complexes in WSe2 (or WS2) monolayers are similar in character to those of MoSe2 for the case of hole doping, but are qualitatively different for the case of electron doping. This is because the sign of the CB spin-orbit splitting is opposite in W-based TMD monolayers in comparison to MoSe2. As shown in Fig. 2, photoexcited electrons are promoted to the upper CBs, rendering the photoexcited e-h pair distinguishable in electron-rich conditions as long as the upper CBs are not populated by resident electrons. The photoexcitation involves distinct resonances for the neutral exciton in Fig. 2(a), singlet trion in (b), triplet trion in (c), and the optimal hexciton complex in (d). The hexciton comprises six particles: the photoexcited e-h pair, and two particle-hole pairs from the two lower CBs at ±$\pm K$ [25, 24, 39] (Appendix A). When the upper CB valleys start to fill at high doping levels, the photoexcited e-h pair can form an eight-particle oxciton by binding to an additional particle-hole pair [24], as shown in Fig. 2(e). The oxciton becomes the optimal complex but it has indistinct resonance at $B$=0. A strong magnetic field can turn this resonance distinct again by emptying the upper CB at $K$, as shown in Fig. 2(f).
Figure 2: Photoexcitation of the $K$ valley in an electron-rich WSe2 monolayer, where optically-allowed transitions couple to the upper CBs. Valleys with solid (dashed) lines host spin-up (spin-down) electrons. (a)-(d) Photoexcitation of charge-neutral excitons ($X^{0}$), negative trions ($X^{-}_{S,T}$), and hexcitons ($H$) when the upper CBs are not populated. (e)-(f) Photoexcitation of oxcitons ($O$) when the upper CBs are populated, without and with a magnetic field. All optical transitions except for (e) are distinct resonances.
Type-B exciton complexes at higher energies
Thus far, we have discussed optical transitions originating from the uppermost VBs at ±$\pm K$, which are the well-known “type-A” excitonic resonances that appear on the low-energy side of the optical spectrum. We emphasize that the type-A resonance with lowest energy (the ground state) always belongs to an optimal excitonic complex. However, excitonic resonances also appear at much higher energies when the photoexcited hole originates from the lower of the spin-orbit-split VBs; these are the “type-B” exciton complexes. Their energies exceed those of type-A resonances by approximately the spin-orbit splitting of the VBs: $\Delta_{v}$similar-to\,$\sim$180 meV in Mo-based TMDs and $\Delta_{v}$similar-to\,$\sim$450 meV in W-based TMDs [46]. Figure 3 shows examples of distinct optical resonances of type-B optimal hexcitons. These complexes emerge in hole-rich TMD monolayers (Mo- and W-based) or in electron-rich MoSe2 monolayers when the upper CBs are unpopulated [25].
Figure 3: Photoexcitation of type-B hexcitons in (a) electron-doped MoSe2, (b) hole-doped MoSe2, and (c) hole-doped WSe2 (or WS2) monolayers. All of these six-particle hexcitons are optimal complexes (because the photoexcited e-h pair couples to all available species of resident carriers), and their optical resonances are distinct (because the e-h pair is photoexcited into empty bands).
[…]
III Classification of excitonic complexes in experiment and their universal trends
Figure 4 shows colormaps of the low-temperature reflectance spectra in electrostatically-doped WSe2 and MoSe2 monolayers, taken from Refs. [20, 23, 22, 21]. Spectral windows that correspond to type-A and type-B optical transitions at zero magnetic field are shown in Fig. 4(a) for WSe2  [20] and in Fig. 4(b) for MoSe2 [23]. Helicity-resolved reflectance spectra of type-A optical transitions are shown in Fig. 4(c) for WSe2 at $B=20\,$T [22] and in Fig. 4(d) for MoSe2 at $B=16\,$T [21]. In all of these cases, distinct resonances of optimal excitonic complexes are indicated with arrows. These (and only these) resonances exhibit a pronounced and sustained energy redshift when the charge density increases. Below, we explain what can be learned from these measurements, and elaborate on universalities of the underlying physics.
III.1 Universal energy redshift of distinct resonances of optimal excitonic complexes
We first focus on the energy redshift of distinct resonances that arise from optimal excitonic complexes. At zero magnetic field, these complexes are type-A hexcitons in electron-rich W-based monolayers and type-B hexcitons in any hole-rich TMD monolayer. Their resonances are marked by arrows in Figs. 4(a) and (b), and their transition diagrams are shown in Figs. 2(d), 3(b), and 3(c). Figure 4(a) shows that the type-A hexciton resonance in WSe2 remains sharp and exhibits a pronounced redshift with increasing electron density, which abruptly ceases only when the gate voltage reaches $\sim$11 V (electron density $\sim$5$\times$1012 cm-2) and electrons begin filling the upper CBs [20]. Beyond this point the photoexcited e-h pair becomes indistinguishable from some carriers in the Fermi sea, and the resonance therefore becomes indistinct. On the other hand, type-B hexcitons in hole-doped conditions remain optimal and their resonances remain distinct even at very large hole densities because $\Delta_{v}\gg E_{F}$, as depicted in Figs. 3(b) and (c). These resonances, labeled by $H_{\mathcal{B}}$ in Figs. 4(a) and (b), exhibit a robust redshift and do not broaden even up to the highest reported doping. We note that type-B resonances are typically broader than those of type-A, due to their very short lifetime [47].
Further important insight is obtained by applying magnetic fields $B$ large enough to fully valley-polarize the Fermi sea of resident carriers. This forces certain optimal $X^{\pm}$ trion resonances, which are otherwise indistinct at $B$=0, to become distinct (because the e-h pair is now photoexcited into empty bands; see Fig. 1(f)). This completely changes how they shift and broaden with increasing doping. Figures 4(c) and (d) show helicity-resolved reflectance spectra in WSe2 and MoSe2, respectively, at strong $B$ [22, 21]. The left and right panels show reflectance of $\sigma_{+}$ and $\sigma_{-}$ circularly polarized light, which probes the $K$ and $-K$ valleys, respectively. To analyze the data we must consider the balance between Zeeman and Fermi energies. Once the monolayer becomes doped, a change of 1 V in gate voltage ($V_{g}$) corresponds to a carrier density change of $\sim$1.5$\times$1012 cm-2 in the WSe2 device [22] and $\sim$2.4$\times$1011 cm-2 in the MoSe2 device [21]. In addition, the signs of the $g$-factors are opposite in the ±$\pm K$ valleys, where $|g_{v}|\approx 6$ for the VB, $|g_{c,2}|\approx 4$ for the CB which takes part in the type-A transition, and $|g_{c,1}|\approx 1$ for the CB which takes part in the type-B transition [11]. These $g$-factors mean that hole-rich monolayers can sustain complete valley polarization at larger carrier densities compared with electron-doped MoSe2 monolayers ($g_{v}>g_{c,2}$), which in turn can sustain complete valley polarization at larger carrier densities compared with electron-doped W-based monolayers ($g_{c,2}>g_{c,1}$).
Starting with MoSe2 in Fig. 4(d), the $X^{\pm}$ resonances in $B$=16 T exhibit a pronounced redshift with increasing doping (in marked contrast to their negligible redshift at $B$=0). However, this redshift persists only as long as the optical transition involves a distinguishable e-h pair. The redshift of $X^{-}$ starts from its emergence when the monolayer becomes electron-doped at $V_{g}\sim$6 V and stops at $V_{g}\sim$10 V (indicated by the white dotted line near $X^{-}$ in Fig. 4(d)). Beyond this point, the Fermi energy exceeds the Zeeman energy and the photoexcited valley starts to fill, and the $X^{-}$ resonance is no longer distinct, marking the onset of its broadening and energy blueshift. Referring to the diagram of Fig. 1(f), the redshift ceases and the broadening is enhanced when the electron density is large enough to begin filling the CB valley at $K$.
For the case of $X^{+}$, whose energy diagram is shown by the inset of Fig. 4(d), the redshift of its distinct resonance persists from its emergence when the monolayer becomes hole-doped at $V_{g}\sim-7$ V, and does not stop even at $V_{g}\sim-16$ V (hole density $\sim$2.2$\times$1012 cm-2). The larger value of $g_{v}$ [48, 49] keeps the photoexcited valley free of resident holes across this entire range of gate voltages.
Similarly, Figure 4(c) shows low-temperature differential reflectance spectra of a WSe2 monolayer when $B=20$ T [22]. The $X^{+}$ trion is optimal and its resonance is distinct between 0 and $-1.7$ V (hole density up to $\sim$2.5$\times$1012 cm-2 in this device), during which its redshift is continuous and the resident holes are fully valley polarized. However, this redshift abruptly ceases when the resonance becomes indistinct at larger $V_{g}$ when holes begin populating Landau levels (LLs) in the $-K$ valley ($V_{g}\lesssim-1.7$ V), which also marks the onset of its enhanced broadening.
Finally, we focus on the electron-doped side of Fig. 4(c), where the gate voltage window of fully valley-polarized electrons is narrowest in W-based monolayers because of the smallness of $g_{c,1}$. The emergence of hexciton resonances ($H$) and the concomitant decay of trions ($X^{-}_{S,T}$) takes place when LLs are filled in the lower CB valleys of both ±$\pm K$ [24]. Similar to the case at $B=0$ in Fig. 4(a), the energy redshift of the type-A hexciton resonance persists only as long as the upper CB valleys are empty (i.e., as long as it remains a distinct resonance). The crossover from hexciton to oxciton occurs when the upper CB valley at $-K$ starts to fill, i.e. when $V_{g}\gtrsim 3.5$ V in this device (electron densities $\gtrsim$5$\times$1012 cm-2). The diagram of this distinct and optimal oxciton is shown in Fig. 2(f). The observed $H\,\rightarrow\,O$ crossover is highlighted in the inset of Fig. 4(c), which magnifies the reflectance signal with helicity $\sigma_{+}$ inside the dashed box at the top left corner (Appendix B).
III.2 Suggested origin of the universal energy redshift
These data show that distinct resonances of optimal excitonic complexes exhibit a universal energy redshift when the charge density increases, and further, that these resonances neither decay nor broaden. The lack of decay and broadening can be reasoned by noting that a distinguishable e-h pair can bind to resident carriers with all available quantum numbers without having to perturb the distribution of resident carriers that are not part of the optimal complex (Appendix A). In Sec. IV, we will discuss why and how shakeup processes broaden the optical resonances in cases that the e-h pair is indistinguishable and/or when the complex is suboptimal.
The universal energy redshift is common to distinct resonances of various optimal complexes. This commonality can be attributed to the small but non-zero difference that exists between band-gap renormalization (BGR) and reduction of the binding energy due to screening by resident carriers (namely, $\Delta E=\Delta E_{g}-\Delta E_{b}$). Increasing the charge density from $n_{1}$ to $n_{2}=n_{1}+\delta n$ is equivalent to increasing the screening through the effective dielectric constant from $\epsilon_{1}$ to $\epsilon_{2}=\epsilon_{1}+\delta\epsilon$. The BGR of a semiconductor is attributed to the Coulomb hole effect [50], reflecting the change in energy needed to excite an electron across the band gap when the Coulomb potential at the immediate vicinity of the electron is changed. Considering a small change, such that $\delta\epsilon$ is small and positive, we get [50, 51, 52, 53]
$\Delta E_{g}$ $=$ $E_{g,2}-E_{g,1}=\lim_{r\rightarrow 0}\left[V(\epsilon_{2},{\bf r})-V(\epsilon_{1},{\bf r})\right]$ (1)
$=$ $\delta\epsilon\,\,V^{\prime}(\epsilon_{1},r=0),$
where $V^{\prime}(\epsilon_{1},r)=(\partial V/\partial\epsilon)\big{|}_{\epsilon=\epsilon_{1}}$. The exciton state is calculated with the same Coulomb potential. Using perturbation theory to quantify the change in exciton binding energy, we get
$\Delta E_{b}=\delta\epsilon\langle\,\,\psi_{1}|V^{\prime}(\epsilon_{1},r)|\psi_{1}\rangle,$ (2)
where $|\psi_{1}\rangle$ is the exciton wave function corresponding to $V(\epsilon_{1},r)$. The total shift of the exciton resonance energy becomes
$\Delta E_{X_{0}}$ $=$ $\Delta E_{g}-\Delta E_{b}$ (3)
$=$ $\delta\epsilon\left\langle\,\,\psi_{1}\left|{\left[V^{\prime}(\epsilon_{1},r=0)-V^{\prime}(\epsilon_{1},r)\right]}\right|\psi_{1}\right\rangle.$
Assuming small size exciton, the term in square brackets is merely the derivative of $r$, and we get
$\Delta E_{X_{0}}\simeq-\delta\epsilon\left\langle\,\,\psi_{1}\left|({\partial^{2}V}/{\partial\epsilon\partial r})\,r\right|\psi_{1}\right\rangle<0.$ (4)
The energy change is negative because the potential $V(\epsilon,r)$ is largely proportional to $1/\epsilon r$, so that $\partial^{2}V/\partial\epsilon\partial r>0$. In other words, the exciton energy redshifts under a small increase in the effective dielectric constant (caused by a small increase in charge density). The energy redshift is expected since the reduction in band-gap energy comes from weakening of the Coulomb potential at $r\rightarrow 0$. On the other hand, the electron-hole binding energy is affected by the change to the Coulomb potential at finite distances and not only at $r\rightarrow 0$. Since the difference between Coulomb potentials with $\epsilon_{1}$ and $\epsilon_{2}$ is largest at $r=0$, the reduction in band-gap energy cannot be completely offset by weaker binding energy, leading to a small overall redshift of the resonance energy when the effective screening increases. Since the energy redshift is a result of increased effective screening, similar behavior is observed when monolayer semiconductors are sandwiched between materials with larger dielectric constants or when semi-metallic thin layers, such as graphite, are placed adjacent to the semiconductor [51, 54].
For simplicity, the derivation above was made for a neutral exciton. Yet, the redshift effect should be similar for complexes with larger number of particles because most of the binding energy comes from the photoexcited pair, where additional particles contribute less and less energy. For example, if we consider a WSe2 monolayer that is encapsulated in hexagonal boron nitride, then the gained energy by forming an exciton is about 170 meV compared with an unbound electron-hole pair [55]. Adding an electron to form a trion gains extra $\sim$30 meV [1], adding the next electron to form a hexciton gains extra $\sim$10 meV, and the next electron to form an oxciton gains extra 3 meV [25, 24]. This behavior is similar to that of energy levels of successive electron shells in atoms but with smaller energy scales.
III.3 Optimality and distinguishability of neutral excitons in moiré heterobilayers
The concepts of optimality and distinguishability are universal and can also be applied to heterostructure systems. Unless we deal with the trivial intrinsic limit (i.e., no resident carriers), charge-neutral excitons in single monolayers cannot be both optimal and distinguishable. This restriction can be alleviated in heterobilayers where in addition to spin and valley, the monolayer in which the particle resides is a relevant quantum number. We illustrate this concept using the reflectance contrast spectrum of WSe2/WS2 moiré heterobilayers, shown in Fig. 5 [26]. The CB and VB alignments are such that electrons are added to the WS2 monolayer when the gate voltage is positive, and holes are added to the WSe2 monolayer when the gate voltage is negative. The low-energy spectral window, shown in the left part of Fig. 5, is governed by absorption processes in the WSe2 monolayer. The high-energy spectral window, shown in the right part, is governed by absorption processes in the WS2 monolayer.
Figure 5: Reflectance contrast spectrum of a WSe2/WS2 heterobilayer at low temperatures where the twist angle between the monolayers is 60. Taken from Ref. [26]. The left and right panels show spectral regions of intralayer excitons in the WSe2 and WS2 monolayers, respectively. The dashed-yellow ovals highlight regions in which the intralayer neutral excitons are optimal and their resonances distinct, supported by their sustained energy redshift and persistent amplitude at elevated charge densities.
To understand how a charge neutral exciton can be optimal and distinguishable in this heterobilayer system, we should exclude the possibility of creating trions with resident carriers. Since electrons reside in the WS2 monolayer, an exciton in the WSe2 monolayer is not affected by these resident electrons except for their screening-induced redshift effect from BGR and binding energy. The reason that a negative trion cannot form is the ineffective interaction between two electrons whose wavefunctions hardly overlap and a hole [45]. As such, the photoexcited e-h pair in the WSe2 monolayer is an optimal exciton, and the lack of resident electrons in this monolayer renders the e-h pair distinguishable. The intralayer neutral exciton in the WSe2 monolayer at positive voltages is therefore an optimal complex with distinct resonance, whose hallmark is sustained energy redshift, as highlighted in the dashed yellow oval around 1.68 eV in Fig. 5. Here, adding resident electrons to the WS2 monolayer screens the Coulomb potential in the WSe2 monolayer, and the ensuing interplay between BGR and binding energy (Sec. III.2) brings in the observed energy redshift. Equivalently, a positive trion cannot form when the neutral exciton resides in the WS2 monolayer and holes reside in the WSe2 monolayer. The result is an optimal intralayer exciton in the WS2 monolayer with distinct resonance, whose sustained energy redshift is highlighted in the dashed yellow oval around 2 eV in Fig. 5.
Figure 6: Photoluminescence (PL) measured from the spectral window of the dark positive trion in a WSe2 monolayer at $B=9$ T when the hole density is (a) 1.4$\times$1011 cm-2 and (b) 1.7$\times$1012 cm-2. The red (black) line corresponds to detection of circularly polarized emission with $\sigma_{+}$ ($\sigma_{-}$) helicity. Taken from Ref. [11]. The highlighted resonances are of dark trions, $X^{+}_{D}(\pm K)$, and their $K_{3}$-phonon replicas, $X^{+}_{D,K_{3}}(\sigma_{\pm})$. (c) and (d) Diagrams of the corresponding emission processes (see text). We have omitted the optically-active upper spin-split CB valleys (see Fig. 2) since optical transitions of dark trions involve the lower CB valleys.
III.4 Ruling out screening by resident carriers as a source of decay of excitonic resonances
We have suggested that screening by resident carriers can explain the steady energy redshift of distinct resonances of optimal excitonic complexes. We wish to examine if this screening plays a significant role in the decay and broadening of other optical resonances. For example, the resonances of trions in MoSe2 at $B$==\,=00, marked by $X^{\pm}$ in Fig. 4(b), represent indistinct resonances of optimal complexes. The same is applicable for the positive trion in WSe2, marked by $X^{+}$ in Fig. 4(a). These resonances behave qualitatively differently than distinct resonances of optimal complexes. When the charge density increases, indistinct resonances of optimal complexes decay and broaden, while showing energy blueshift at elevated charge densities. We will first address their decay and broadening and in Section IV we will focus on their energy shift, which goes beyond the considerations of band filling, BGR, and binding energy.
To better understand the origin for decay and broadening of indistinct resonances, we compare the behavior of $X^{+}$ to that of $H_{\mathcal{B}}$ either in Figs. 4(a) or (b). While type-A positive trions are quenched when the gate voltage is approaching  $-10\,$V, the amplitudes and linewidths of the type-B hexcitons are not affected by increasing the density of holes. Since charge particles screen charge particles, the screening effect cannot be selective in weakening the Coulomb attraction between the particles of one complex species but not of another. Thus, the different behavior of $X^{+}$ and $H_{\mathcal{B}}$ in either Figs. 4(a) or (b) suggests that screening does not play a significant role in the decay and broadening of indistinct resonances of optimal excitonic complexes (at least in the studied range of charge densities of these experiments). That distinct resonances of optimal excitonic complexes neither decay nor broaden at the same charge density that other resonances do broaden, means that the distinguishability of the photoexcited e-h pair and optimality of the excitonic complex play the important roles in setting their decay and broadening.
Clear-cut evidence against screening as a source of decay and broadening can be inferred by studying the recombination of dark trions in tungsten-based monolayers. Figure 6 shows the photoluminescence (PL) of dark positive trions in a WSe2 monolayer at $B=9$ T when the hole density is (a) 1.4$\times$1011 cm-2 and (b) 1.7$\times$1012 cm-2 [11]. We first explain what is being measured and then we relate the findings to the role of screening. The red (black) curve corresponds to detection of circularly polarized emission with $\sigma_{+}$ ($\sigma_{-}$) helicity. Emission due to recombination of dark trions without phonon assistance is denoted by $X^{+}_{D}(-K)$ and $X^{+}_{D}(K)$ in Figs. 6(a) and (b), and the corresponding recombination diagrams are shown in Figs. 6(ci) and (di). The unpolarized light emission is a result of the out-of-plane dipole orientation [56], yielding similar amplitudes for emission with $\sigma_{+}$ and $\sigma_{-}$ helicity. Polarized emission is restored when the recombination is assisted by zone-edge phonons, such as the $K_{3}$ mode, through which the electron experiences an intervalley transition and then recombines with a hole of matching spin in the opposite valley [57, 18]. This emission is denoted by $X^{+}_{D,K_{3}}(\sigma_{\pm})$ in Figs. 6(a) and (b), and the corresponding recombination diagrams are shown in Figs. 6(cii) and (dii).
Coming back to the screening question, we first notice that we are dealing with different recombination processes of the same optimal trion. For example, we say that the optical transition in Fig. 6(ci) is distinct because the recombining hole belongs to an empty valley (VB valley at $-K$). On the other hand, we can also say that the phonon-assisted optical transition in Fig. 6(cii) of the very same optimal trion is indistinct because now the recombining hole belongs to a populated valley (VB valley at $K$). One cannot argue that broadening and decay of $X^{+}_{D,K_{3}}(\sigma_{+})$ in Fig. 6(b) compared with Fig. 6(a) is caused by screening because the very same trion shows no signs of decay and broadening when measured through its emission from $X^{+}_{D}(-K)$. In fact, the latter resonance is stronger and slightly narrower at larger charge density.
Similarly, we can look at dark positive trions whose electron resides in the opposite valley, as shown in Figs. 6(di) and  6(dii). As before, resonances of this optimal trion are said to be distinct or indistinct based on the identity of the recombining hole. Inspecting the resonance $X^{+}_{D,K_{3}}(\sigma_{-})$ in Figs. 6(a) and (b), we see that increasing the hole density renders the phonon-assisted emission narrower and stronger, corroborated by the fact that this is a distinct resonance. On the other hand, the indistinct resonance $X^{+}_{D}(K)$ in Figs. 6(a) and (b) shows significant decay and broadening when the charge density increases. We therefore cannot conclude that screening is causing this decay, because emission from the very same trion is narrower and stronger if it involves recombination of the distinguishable hole.
Figure 7: Helicity resolved magneto-optical reflectance spectra of the neutral exciton in a charge tunable WSe2 monolayer at 4 K. Taken from Ref. [7]. The out-of-plane magnetic field is 17.5 T. As a guide to the eye, the faint dashed lines trace the energy blueshift of the exciton. (b) Helicity resolved optical transitions when the electrons in the monolayer are fully valley polarized. (c) Helicity resolved optical transitions when the holes are fully valley polarized.
III.5 The secondary role of band filling as a source of energy blueshift in absorption
We have identified in Sec. III.1 that distinct resonances of optimal excitonic complexes redshift in energy when resident carriers are added to the system. However, when indistinguishable e-h pairs are photoexcited, absorption resonances trend to blueshift. One should therefore account for the expected energy blueshift from Pauli blocking (i.e., band filling) – namely, the optical energy gap increases because resident carriers occupy the low energy states in the photoexcited valley up to the Fermi energy. Using WSe2 to compare the energy shifts seen in electron and hole doped monolayers, it will become clear that band filling is not the primary reason for the observed energy blueshift of indistinct resonances.
Figure 7(a) shows low-temperature differential reflectance spectra of the neutral exciton in a WSe2 monolayer when $B$=17.5 T [7]. Changing the gate voltage by 1 V in this device amounts to a change of $\sim$4$\times$1011 cm-2 in charge density. Given the large magnetic field, the electrons are fully polarized in the lower CB valley at $-K$ when applying a small positive gate voltage, as shown in Fig. 7(b). Similarly, the holes are fully polarized in the upper VB valley at $K$ when applying a small negative gate voltage, as shown in Fig. 7(c). The gate-voltage window of complete valley polarization is larger for holes because of the larger $g$-factor of the VB valley [11].
We see that the energy blueshift of the exciton is more than 20 meV when the gate voltage increases from 0 to 4 V, and that the blueshift is independent of the helicity (i.e., similar for $\sigma_{\pm}$ polarized light). As we can infer from Fig. 7(b), photoexcitation of either valley can create suboptimal neutral excitons with distinct resonances. The situation is qualitatively different in a fully valley-polarized WSe2 monolayer that is electrostatically-doped with holes. As shown in Fig. 7(c), photoexcitation of the valley at $K$ creates optimal neutral excitons with indistinct resonances. Examining the behavior in Fig. 7(a) when $V_{g}<0$, the energy blueshift of the optimal exciton from $K$ has a staircase shape. As indicated by the optical transition with $\sigma_{+}$ polarization in Fig. 7(c), the staircase shape is governed by progressively blocking the optical transition due to continuous filling of Landau levels with holes. Contrary to this case, the optical transition with $\sigma_{-}$ polarization in Fig. 7(c) means that a suboptimal exciton from $-K$ is not subjected to band filling (i.e., no Pauli blocking). And yet, Fig. 7(a) shows that the energy blueshift of the distinct resonance of the suboptimal exciton from $-K$ is more than 20 meV when the voltage changes from 0 to nearly -4 V whereas the indistinct resonance of the optimal exciton from $K$ experiences similar blueshift when the voltage changes from 0 to -8 V. Namely, the energy blueshift of the neutral exciton is doubled when an empty valley with no Pauli blocking is photoexcited compared with the blueshift that occurs when photoexcitation takes place in an occupied valley (i.e., when Pauli blocking should be strongest). As discussed below, this surprising result can be understood by considering shakeup processes.
IV Shakeup Processes
We have seen that screening is not a primary source of decay, and that band filling is not a major source of energy blueshift in TMD monolayers. These seemingly counterintuitive phenomena can be settled by recognizing that screening by charge carriers and band filling effects are mostly significant when the Fermi energy is comparable to or larger than the binding energy of the e-h pair [50, 58, 59]. In the opposite limit ($\varepsilon_{X}\gg E_{F}$), applicable to TMD systems, we argue that broadening and energy shift of excitonic complexes are instead primarily governed by shakeup processes of resident carriers. In general, there are two types of shakeup processes. The first one involves a breakup (i.e., physical separation) of distinguishable and indistinguishable resident carriers upon photoexcitation (or their reunion after recombination), applicable in the case of indistinct resonances of optimal complexes. The second process involves the evacuation (i.e., removal) of distinguishable resident carriers from the creation region of suboptimal excitonic complexes. We elaborate on these processes in this section.
Figure 8: Colormaps of (a) reflectance and (b) photoluminescence spectra in MoSe2 at $B=0$. Taken from Ref. [23]. The indistinct resonances of the optimal trion complexes, labeled by $X^{\pm}$, broaden when resident carriers are added to the monolayer (readily seen when $|V_{g}|\gtrsim 4$ V). (c) Diagram of the positive trion following photoexcitation of the $K$ valley. (d) Breakup process following the creation of $X^{+}$. (e) Reunion process following the recombination of $X^{+}$. These processes correspond to the broadening and energy blueshift (redshift) in absorption (emission), seen inside the highlighted dashed ovals in (a) and (b).
IV.1 Energy shifts and broadening of indistinct resonances of optimal complexes through breakup and reunion processes
When distinguishable and indistinguishable resident carriers are present, breakup or reunion processes govern the behavior of indistinct resonances of optimal complexes. We first explain these processes in charge tunable MoSe2 monolayers at $B$=0, due to its relatively simple absorption and emission spectra, shown in Figs. 8(a) and (b) [23]. In absorption (emission), the indistinct resonances of optimal trions blueshift (redshift) and broaden in energy. Without loss of generality, we analyze these phenomena by focusing on the indistinct resonance $X^{+}$ whose absorption process in the $K$ valley is shown in Fig. 8(c).
Prior to the absorption of $X^{+}$, resident holes from the VB valleys at $-K$ and $K$ have some spatial overlap. This overlap is illustrated through the wave-packet envelopes of the holes in the left diagram of Fig. 8(d). Mandated by the Pauli exclusion principle, this overlap breaks when the resident hole from $-K$ binds to the photoexcited e-h pair from $K$ to form $X^{+}$, illustrated by the right diagram of Fig. 8(d), where the arrow signifies that resident holes from $K$ are expelled from the region of the created trion. The energy of the absorbed photon is converted to $\hbar\omega=\varepsilon_{T}+\delta_{\varepsilon}$, where $\varepsilon_{T}$ is the trion energy and $\delta_{\varepsilon}$ is the energy required to break the overlap, through which we can understand the energy blueshift. Since the overlap is larger with increasing charge density and since it costs more energy to break a stronger overlap, the energy blueshift is scaled with charge density. Furthermore, the energy blueshift of $X^{+}$ in Fig. 8(a) indicates that $\delta_{\varepsilon}$ is stronger than the underlying redshift effect from screening at elevated densities (Sec. III.2). Finally, the broadening of $X^{+}$ stems from the short lifetime of the perturbed resident-carriers system, and because $\delta_{\varepsilon}$ has a continuum of energy costs depending on whether the breakup involves emission of long wavelength plasmons (with various wavevectors) or trion-hole exchange scattering [18].
Figure 8(e) shows the respective process during recombination. Mandated by the Pauli exclusion principle, the left diagram shows the lack of overlap between the wave-packet envelopes of a hole from $K$ and a trion with indistinguishable e-h pair from $K$. After emission, the overlap between the left-behind hole from $-K$ and a resident hole from $K$ is restored, as shown by their reunion in the right diagram of Fig. 8(e). The recombination process leaves behind an inhomogeneous Fermi sea because the region that was occupied by the trion is left void of indistinguishable resident carriers, which are then pushed back to replenish this region as indicated by the arrow in Fig. 8(e). The energy of the trion before recombination is converted to $\varepsilon_{T}=\hbar\omega+\delta_{\varepsilon}$, where $\delta_{\varepsilon}$ in this case is the energy difference between the inhomogeneous Fermi sea immediately after recombination and the homogeneous one in equilibrium. Since $\delta_{\varepsilon}$ is commensurate with the hole density, the conversion leaves the emitted photon with less energy when the charge density increases, in line with the measured energy redshift inside the highlighted dashed white oval in Fig. 8(b). The redshift is further enhanced by the interplay between BGR and reduced binding energy due to screening (Sec. III.2). As before, the broadening reflects the short lifetime of the perturbed system. That is, the broadening is a measure of the time it takes the inhomogeneous Fermi sea immediately after recombination to restore homogeneity. This response time is commensurate with $\omega_{p}^{-1}$, where $\omega_{p}$ is the plasma frequency.
The self-consistency of this physical picture is examined through the unique case of WSe2 monolayers. Figure 9(a) shows the spectral window of trion emission in a charge tunable WSe2 monolayer at T $=$ 4 K. Details of the sample fabrication and experimental setup can be found in the supplementary information of Ref. [60], and analysis of the dark exciton and trion resonances, $D^{0},\,D^{+},\,D^{-},\,D^{-}_{B}$ can be found in Ref. [18]. Figures 9(b) and (c) show the extracted resonance energies and their full width at half maximum (FWHM), respectively, where the dashed lines show the resonances energies at T $=$ 30 K as well. The numbers in Fig. 9(b) denote the energy redshift of the corresponding resonance in meV per 1011 cm-2. The energy redshift of $H$ and its lack of broadening, as summarized in Figs. 9(b) and (c), are hallmarks of the distinct resonance of the optimal hexciton. The resonance does not broaden since the Fermi sea is not perturbed in this case, and as discussed in Sec. III.2, its energy redshift stems from the interplay between BGR and reduced binding energy due to screening.
Figure 9: Photoluminescence intensity colormap of WSe2 at $B\,$==\,=0, shown in the spectral window of trions. The laser energy is 1.96 eV and its power is 1 $\mu$W. (b) and (c) Extracted peak energies and full-width at half maximum of the resonances in (a). The numbers in (b) denote the energy redshift in meV per 1011 cm-2, where a change of 1 V in this device corresponds to a change of 1011 cm-2 in charge density.
The energy redshift of $X^{+}$ in both MoSe2 and WSe2, evidenced from the emission spectra in Figs. 8(b) and  9(a), comes from the fact that their positive trions are optimal complexes with indistinct resonances. Namely, the emission of $X^{+}$ is accompanied by a reunion process. The enhanced broadening of $X^{\pm}$ in Fig. 8(b) compared with that of $X^{+}$ in Fig. 9(a) is attributed to the larger densities used in the MoSe2 device, where --10 V in Fig. 9(a) is the equivalent of --1.5 V in Fig. 8(b). Because recombination of an optimal dark trion at $B=0$ is governed by optical transition of an indistinguishable e-h pair, reunion processes are also the reason that dark trion resonances broaden and redshift in energy, as shown in Fig. 9. Finally, of all resonances in the emission spectra of Figs. 8(b) and  9(a), the only ones that do not shift in energy are the triplet and singlet negative trions ($X^{-}_{S,T}$). As we show next, this behavior is consistent with the fact that $X^{-}_{S,T}$ are distinct rather than indistinct resonances.
The energy diagrams of the triplet, singlet, dark positive trion, and hexciton in WSe2 monolayer are illustrated in parts (i) of Figs. 10(a)-(d). The use of different colors is meant to distinguish between particles with different spin and valley quantum numbers. As shown in Fig. 10(a), there is no restriction for an electron from the bottom CB valley at $K$ to be completely excluded from the immediate vicinity of the triplet trion, illustrated by the slight overlap of $X^{-}_{T}$ and $e_{K}$ in diagram (ii). The same is applicable for an electron from the bottom CB valley at $-K$ and the singlet trion as shown by the corresponding diagram in Fig. 10(b). After recombination of the trion, the electron that is left behind is already in proximity with an electron from the time-reversed valley, as shown by diagrams (iii) of Figs. 10(a) and (b). That is, the distribution of resident carriers remains homogeneous in the presence of $X^{-}_{S,T}$. The result is that the distinct resonances of $X^{-}_{S,T}$ do not redshift in energy because there is no need for reunion processes to follow the recombination. Contrary to these cases, other trion species require reunion processes because the Pauli exclusion principle is at play, as illustrated in Fig. 10(c) for the positive dark trion. Finally, Fig. 9(a) shows that when the electron density is larger than 1012 cm-2 ($V_{g}>10\,$V), $X^{-}_{S,T}$ decay and $H$ emerges. The corresponding hexciton diagrams are shown in Fig. 10(d), where the e-h pair simultaneously binds to two nearby electrons from the bottom CB valleys at $K$ and $-K$.
Figure 10: Diagrams of the triplet and singlet negative trions, positive dark trion and hexciton in WSe2 monolayer. For each complex, (i) corresponds to its energy diagram, and (ii)/(iii) is the corresponding real-space configuration before/after recombination. Of the 4 shown complexes, only $D^{+}$ involves recombination of an indistinguishable e-h pair, and therefore its emission is followed by a reunion process.
Comparing the behavior of $X^{-}_{S,T}$ to that of $H$, we should explain why $X^{-}_{S,T}$ are not subjected to the same screening-induced energy redshift like $H$, and what causes the broadening of $X^{-}_{S,T}$. The broadening indicates a shorter lifetime when the electron density increases, caused by the trion-electron interaction which renders them dark through exchange scattering (i.e., $X^{-}_{S,T}\rightarrow D^{-}$) [18]. The lack of screening-induced energy redshift, on the other hand, is less clear. We can only conjecture that if the resident electrons are quasi-localized when the electrons density $n_{e}$ is a few times $10^{11}$ cm-2 or less, then $X^{-}_{S,T}$ in Fig. 10 are not subjected to the screening-induced energy redshift that $H$ is subjected to when the resident electrons become itinerant at $n_{e}>10^{12}$ cm-2 (Appendix D).
Finally, the emission spectra in Figs. 8(b) and  9(a) show that the energy redshift of indistinct resonances of optimal complexes exist at all charge densities, and that their redshift and broadening are enhanced at elevated charge densities. The enhancement in the case of $X^{\pm}$ in Fig. 8(b) is evident when $|V_{g}|\gtrsim 4\,$$V$, and in the case of $D^{-}$ and $D^{-}_{B}$ in Fig. 9(b) when $V_{g}\gtrsim 15\,$$V$. Our suggested explanation is that reunion processes are relevant regardless of charge density, where the extent of the energy redshift and broadening depend on whether the resident carriers are quasi-localized or itinerant. In more detail, since the Pauli exclusion principle precludes overlap of trions with indistinguishable resident carriers, the rate at which indistinguishable resident carriers creep into the trion region after recombination depends on their itinerancy. The smaller rate of energy redshift and broadening when the resident carriers are quasi-localized is consistent with the fact that Coulomb correlations help indistinguishable and distinguishable resident carriers to avoid each other better at small charge densities [61]. The resulting overlap between indistinguishable and distinguishable resident carriers is smaller than the one expected if they were itinerant, leading to mitigated redshift and broadening effects.
IV.2 Energy blueshift of neutral excitons through evacuation of distinguishable resident carriers
Neutral excitons in TMD monolayers have three possible configurations: a suboptimal exciton with distinct resonance (e.g., Fig. 2(a)), a suboptimal exciton with indistinct resonance (e.g., Fig. 1(a)), or an optimal exciton with indistinct resonance (e.g., Fig. 1(e)). As we discussed in Sec. III.5, while suboptimal excitons with distinct resonances are not subjected to band filling effects (Pauli blocking), they show evidently stronger energy blueshift in absorption experiments compared with the indistinct resonance of the optimal exciton which is subjected to the strongest band filling effect (Fig. 7).
Figure 11: (a) Absorption diagram of a suboptimal neutral exciton with indistinct resonance in electron-doped MoSe2 monolayer. (b)/(c) Corresponding real-space configuration before/after absorption. The bolder arrow of the distinguishable resident carrier indicates a stronger shakeup.
To elucidate the origin of this behavior, we should understand the different roles played by distinguishable and indistinguishable resident carriers. Figure 11(a) shows the photoexcitation of a suboptimal neutral exciton with indistinct resonance, wherein half of the resident electrons are distinguishable ($-K$ valley) and the other half are indistinguishable ($K$ valley). Figures 11(b) and (c) show the conditions before and after the photoexcitation, where the boldness and length of the arrows in Fig. 11(c) reflect the perturbation to distributions of distinguishable and indistinguishable resident carriers. Since the Pauli exclusion principle prohibits the creation of the exciton in the immediate vicinity of indistinguishable resident electrons, it costs more energy to evacuate nearby distinguishable resident carriers with which the e-h pair can bind than to rearrange the more remote indistinguishable resident carriers with which the pair cannot bind [45, 61].
Figure 12: (a) Energy shift and FWHM of the neutral exciton resonance as a function of charge density, extracted from low-temperature reflectance spectra of hBN-encapsulated WSe2 monolayer. (b) The absorption processes in the hole- and electron-doped regimes are shown in the left and right diagrams, respectively. (c)-(e) Corresponding helicity-resolved results in hole-doped monolayer at $B=9\,$T. Also shown are results of the distinct resonance of the optimal positive trion, labeled by $X^{+}(\sigma_{-})$. The energy blueshift and FWHM of $X^{0}(\sigma_{-})$ are shown in a range of hole densities at which this resonance is still identifiable in the reflectance spectra (its oscillator strength is rapidly depleted in favor of $X^{+}(\sigma_{-})$).
We demonstrate how this physics manifest through the absorption resonances of neutral excitons in WSe2 monolayers. The choice to focus on absorption resonances in this monolayer is intentional because we can then readily decipher the roles of distinguishability and optimality in all three possible configurations of the exciton under various conditions. On the other hand, the recombination and energy relaxation of neutral excitons involve exchange scattering processes that complicate the analysis of emission resonances [18]. Exchange scattering is also relevant when dealing with absorption resonances of excited-state neutral excitons, and we leave the analysis of this case to Appendix E.
Figure (12)(a) shows the energy shift and FWHM of the neutral exciton absorption resonance as a function of charge density in an hBN-encapsulated WSe2 monolayer at zero magnetic field. These results were extracted from low-temperature reflectance spectra, where details of the sample fabrication and experimental setup can be found in the supplementary information of Ref. [60]. When the holes density increases from zero to 1012 cm-2, the energy blueshift is 7.5 meV and the FWHM increases from 2 to 6 meV. Half of the resident holes are distinguishable and the other half are indistinguishable, as shown by the left diagram of Fig. (12)(b). When the monolayer is electron-doped, the corresponding energy blueshift is 10 meV and the FWHM increases from 2 to 7 meV. The resident electrons are 100% distinguishable in this case, as shown by the right diagram of Fig. (12)(b). Assuming that distinguishability of the resident carriers is the key point regardless of whether these are electrons or holes (as long as their effective masses are comparable), we can use the results of Fig. (12)(a) to infer the following. Increasing the density of distinguishable (indistinguishable) resident carriers by 1011 cm-2 roughly adds 1 meV (0.5 meV) to the blueshift and 0.5 meV (0.3 meV) to the broadening of the neutral exciton resonance. That is, the effect from distinguishable resident carrier is about as twice as strong.
The stronger effect caused by the presence of distinguishable resident carriers is corroborated by helicity-resolved measurements of the energy shift and FWHM from the same device. Figures 12(c) and (d) show the respective results when $B=9\,$T, where diagrams of the corresponding absorption resonances are shown in Figs. (12)(e). We can confirm that the holes are fully valley polarized through the steady energy redshift and constant FWHM of the distinct resonance of the positive trion, labeled by $X^{+}(\sigma_{-})$. Focusing on the neutral exciton diagrams in Fig. 12(e), the valley-polarized resident holes are 100% distinguishable when dealing with the distinct resonance of the suboptimal exciton, labeled by $X^{0}(\sigma_{-})$, and 100% indistinguishable when dealing with the indistinct resonance of the optimal exciton, labeled by $X^{0}(\sigma_{+})$. Comparing the energy blueshifts of $X^{0}(\sigma_{\pm})$ in Fig. (12)(c), we again notice that distinguishable resident carriers induce as twice as strong energy blueshift. The FWHM results in Fig. (12)(d) show that the broadening rate of $X^{0}(\sigma_{+})$ is strongly suppressed compared with that of $X^{0}(\sigma_{-})$. Given that there are no distinguishable resident carriers with which an optimal exciton can form a trion, the broadening of $X^{0}(\sigma_{+})$ is not limited by the same ultrafast process that converts suboptimal excitons to optimal trions (i.e., $X^{0}(\sigma_{-})\,\rightarrow X^{+}(\sigma_{-})$).
Finally, to be consistent with the analysis in Sec. IV.1, we also explain how the energy blueshift of $X^{0}$ emerges in case that the resident carriers are localized. When the density of distinguishable resident carriers increases, the energy of the exciton has to increase as well (i.e., to blueshift) for its state to remain extended, or else the exciton binds to a localized electron and becomes a localized trion [62, 61]. When the density of indistinguishable resident carriers increases, the energy of the exciton has to increase, albeit at a lesser blueshift rate, because of the smaller region left in the sample for the exciton to form [61].
V Compressibility and exciton resonances
In previous sections, we have implicitly assumed that the two-dimensional gas of mobile carriers is compressible, meaning that the resident carriers can scatter in response to absorption or emission of the e-h pair (e.g., to accommodate the Pauli exclusion principle). This response is impeded in TMD heterostructures with fractionally filled moiré lattices or in semiconductors with LLs at integer fillings wherein the macroscopic state of the resident carriers is incompressible. That is, when the system enters an insulating ground state with a finite energy gap below the excited state. In the following, we first present relevant experimental results and then discuss how compressibility of the resident carrier gas is related to the amplitude and energy shift of excitonic resonances.
Starting with a monolayer system, Fig. 13 shows $B$-dependent polarized optical absorption spectra of hBN-encapsulated WSe2 [8] (we note that these colormaps show $B$-dependent spectra at fixed carrier density, in contrast to density-dependent spectra at fixed $B$ shown in previous figures). The hole density in Figs. 13(a) and (b) is 1.7×$\times 10^{12}$ cm-2. The emergence of exciton resonances in Fig. 13(a) and the onset of energy blueshift of $X^{+}$ in Fig. 13(b) take place when $B\gtrsim 10$ T. The hole density in Figs. 13(c) and (d) is 4.6×$\times 10^{12}$ cm-2, for which a similar behavior is observed when $B\gtrsim 25$ T. These strong magnetic fields are needed to fully valley-polarize the monolayer where holes only populate LLs of the VB valley at $K$, as shown in Fig. 13(e). Photoexcitation of this valley creates optimal excitons with indistinct resonance ($\sigma_{+}\rightarrow X^{0}$), while photoexcitation of the valley at $-K$ creates optimal positive trions with distinct resonance ($\sigma_{-}\rightarrow X^{+}$).
The energy blueshift of $X^{+}$ when $B$ increases in Figs. 13(b) and (d) can be understood from the larger $g$-factors of the VB [11], where the Zeeman energy lowers the VB valley at $-K$ more than it lowers the CB valleys at $-K$ (see Fig. 13(e)). The dashed white lines in Figs. 13(a) and (c) trace the energy redshift from the optical transition in $K$. They have opposite slope to the energy blueshift in Figs. 13(b) and (d) due to the opposite-sign $g$ factors of the opposite valleys, $g_{-K}=-g_{K}$. That is, the Zeeman energy raises the VB valley at $K$ more than it raises the CB valleys at $K$. The dashed white lines in this case trace the energy redshift of $X^{0}$ at exact integer filling of the LLs, where the labeled integers indicate the highest filled LL of the VB valley at $K$. The indistinct resonances of the optimal exciton in Fig. 13(c) appear above 1.76 eV whereas those in Fig. 13(a) appear below 1.74 eV due to different numbers of filled LLs (i.e., weaker/stronger role of Pauli blocking at smaller/larger hole densities).
The unique behavior we wish to focus on is the repeated blueshift and amplitude modulation of the resonance $X^{0}$ between exact integer filling of the LLs, shown when the exciton resonance departs from the white dashed line and extends into the dotted darker lines in Figs. 13(a) and (c). We glean two important insights from these measurements.
  1. 1.
    The absorption amplitude of $X^{0}$ nearly vanishes at integer LL fillings and reaches a maximum at half fillings. This behavior can be seen by comparing the amplitudes along the dashed white line and those midway through the dotted black lines in Figs. 13(a) and (c). In contrast, Figs. 13(b) and (d) show that the absorption amplitude of $X^{+}$ is hardly affected by the filling factor, consistent with the fact that the photoexcited pair is distinguishable and holes are fully polarized. That is, because $X^{+}$ is a distinct resonance of an optimal positive trion.
  2. 2.
    When the $(\ell+1)$ LL is partially filled, the energy blueshift of the exciton resonance along a black dotted line is qualitatively described by
    $E_{X^{0}}(B)\approx E_{\ell+1}+C\times\frac{B-B_{\ell+1}}{B_{\ell}-B_{\ell+1}}\,,$ (5)
    where $B_{\ell}\geq B\geq B_{\ell+1}$. $B_{\ell}$ and $B_{\ell+1}$ are the magnetic fields needed to fill exactly $\ell$ and $(\ell+1)$ LLs, respectively. $E_{\ell+1}$ is the exciton resonance energy when exactly ($\ell+1$) LLs are filled, and $C$ is an energy constant that depends on charge density.
Figure 13: Colormaps of helicity resolved magneto-optical absorption spectra of an electrostatically hole-doped WSe2 monolayer as a function of magnetic field and photon energy at 4 K. Taken from Ref. [8]. The measured signal is $(1-T/T_{0})^{\pm}$, where $T$ ($T_{0}$) is the transmission (reference) spectrum, and $\pm$ refers to light with $\sigma_{\pm}$ polarization. The hole density is 1.7×$\times 10^{12}$ cm-2 in (a)-(b) and 4.6×$\times 10^{12}$ cm-2 in (c)-(d). (e) Schemes of optical excitations in valley-polarized WSe2 monolayer with hole Landau levels at the VB valley of $K$.
Figure 14: (a) Schemes of interlayer exciton complexes within the moiré unit cell of WS2/WSe2 heterobilayers. (b) Scheme of the moiré superlattice when the fractional hole filling is $\nu=-1/3$. (c) and (d) Low temperature PL intensity plots of the interlayer exciton versus gate voltages in R- and H-stacked devices, respectively. Taken from Ref. [27]. The corresponding optical excitations are 1.678 and 1.96 eV, and the laser powers are 50 nW. These measurements were taken at zero magnetic field.
A behavior with similar trends is observed in PL experiments of WS2/WSe2 heterobilayers performed by Wang et al. [27]. Figure 14(a) shows schemes of the interlayer exciton complex in the moiré unit cell. The top scheme corresponds to R-type stacking in which the electron and hole are aligned. The bottommost CB of the heterobilayer belongs to the top monolayer (WS2), and the topmost VB belongs to the bottom monolayer (WSe2). The bottom scheme corresponds to H-stacking (60 twist angle between WS2 and WSe2), wherein the lowermost part of the CB in the top monolayer is spread across the moiré unit cell [27]. Figure 14(b) shows a scheme of the moiré superlattice when the fractional hole filling is $\nu=-1/3$. The highlighted parallelogram encompasses the area of three moiré unit cells, within which the total charge is one hole.
The rich PL behavior of the interlayer moiré excitons is shown in Figs. 14(c) and (d) as a function of gate voltage and photon energy for R- and H-stacked devices, respectively [27]. The horizontal white dashed lines correspond to gate voltages with integer filling of the moiré unit cells. The gate voltage at charge neutrality is marked by $\nu=0$. At larger (smaller) gate voltages, electrons (holes) are added to the WS2 (WSe2) monolayer and $\nu>0$ ($\nu<0$). The black dotted lines in Figs. 14(c) and (d) trace the energy blueshift of the interlayer moiré exciton when electrons or holes are added to the monolayer. The left edges of these black dotted lines mark voltages of exact fractional filling in the moiré superlattice. The PL behavior of the interlayer exciton in Fig. 14 when changing the charge density is not foreign to the absorption behavior of $X^{0}$ in Fig. 13 when changing the magnetic field. Namely,
  1. 1.
    The PL intensity of the interlayer moiré exciton nearly vanishes at exact fractional filling, and reaches maximum intensity when the charge densities are between two fractional states (mid section of the black dotted lines).
  2. 2.
    The energy blueshift of the interlayer moiré exciton along the black dotted lines in Figs. 14(c) and (d) can be described as
    $E_{IX}(n)\approx E_{\nu^{\prime}}+\mathcal{C}\times\frac{n-n_{\nu^{\prime}}}{n_{\nu}-n_{\nu^{\prime}}}\,,$ (6)
    where $n_{\nu}\geq n\geq n_{\nu^{\prime}}$. $n$ is the charge density, limited between the charge densities of adjacent fractionally-filled states, $n_{\nu}$ and $n_{\nu^{\prime}}$ ($|\nu|>|\nu^{\prime}|$). Since charge density is a positive quantity, we use assignments such as $\{\nu=1/3,\,\nu^{\prime}=1/4\}$ in the electron-doped case versus $\{\nu=-1/3,\,\nu^{\prime}=-1/4\}$ in the hole-doped case. $E_{\nu^{\prime}}$ is the resonance energy at fractional filling $\nu^{\prime}$ (left edges of the black dotted lines), and $\mathcal{C}$ is an energy constant that depends on the charging energy of the moiré unit cell.
Suggested interpretation
The experimental findings show that the absorption amplitude of the neutral exciton vanishes at exact integer filling of LLs in a TMD monolayer, and that the emission amplitude of the interlayer exciton vanishes at exact fractional filling of the TMD moiré heterobilayer. Both cases suggest that the exciton cannot exist if the state of the resident carriers is incompressible. To accommodate the presence of the exciton in a monolayer semiconductor, indistinguishable resident carriers must be able to elastically scatter out of the area occupied by the exciton or else the Pauli exclusion principle is violated [7, 61]. We emphasize that the presence of resident carriers is not necessary for emission or absorption of an exciton; these processes are strongest at charge neutrality without any resident carriers. The argument is that if resident carriers are present in the system, then they must reside in a compressible state for an exciton to exist.
The elastic scattering in the WS2/WSe2 heterobilayer system is needed because the repulsion between overlapping distinguishable electrons in the WS2 monolayer is stronger than their attraction to a hole in the WSe2 monolayer. In other words, the interlayer exciton is formed when one of the two electrons scatters out of the exciton region. Such scattering is feasible with compressible states (e.g., a non-fractional state in the moiré heterobilayer), where the maximum PL intensity occurs midway between two fractional states because elastic scattering is strongest at these conditions. That is, when $f(1-f)=1/4$ where $f=1/2$ and $(1-f)=1/2$ are the occupation probability of a resident carrier prior and after scattering, respectively. The resident carriers can then scatter relatively freely and enable the shakeup processes.
Similarly, exciton formation is inhibited in a monolayer subjected to LL quantization at integer filling. Without loss of generality, we explain this physics in a hole-doped monolayer (e.g., Figs. 13(a) and (c)). The hole of the exciton is not subjected to LL quantized motion because the exciton is charge neutral. At integer filling conditions, the holes in the LLs cannot make room for the indistinguishable hole of the exciton (with similar spin and valley). Having a partially filled LL alleviates this constraint because holes from this LL can hop between cyclotron orbit centers. That is, indistinguishable resident holes avoid the exciton by elastic scattering away from the photoexcited hole.
Finally, we notice that the energy blueshift of $X^{0}$ along the dotted black lines in Figs. 13(a),  13(c),  14(c) and  14(d) are all related to the density of compressible resident carriers. This dependence is implicit in the case of a partially filled LL, denoted by the fraction term in Eq. (5), where the density of resident carriers in a LL is proportionate to the magnetic field. This dependence in explicit in the moiré heterobilayer system, evidenced from the fraction term in Eq. (6), which denotes the relative density of itinerant resident carriers between adjacent incompressible states. Deriving the direct relation between compressibility and energy shift, through which one can better quantify the phenomenological $C$ terms in Eqs. (5)-(6), will be studied elsewhere.
VI conclusion
The systematic analysis presented in this work shows that distinguishability and optimality, particularly as they relate to the Pauli exclusion principle, are key factors in determining the energy shift and broadening of various excitonic resonances in electrostatically-doped semiconductor systems. We have corrected common misconceptions, showing that screening by charge particles and band filling are in fact minor reasons for the broadening or energy blueshift of excitonic resonances. The one manifestation of screening is the energy redshift of distinct resonances of optimal excitonic complexes, where Pauli exclusion plays minimal role. The screening leads to a reduction of the bandgap energy and a reduction of the binding energy, where the former is slightly larger than the latter, leading to a small net energy redshift when the charge density increases. This effect is seen through distinct resonances of optimal complexes since their excitation (or recombination) does not involve a shakeup of the resident carrier distribution, and thus, these resonances neither broaden nor decay.
Shakeup processes do, however, govern the energy shifts and broadening of indistinct resonances of optimal excitonic complexes. In absorption, these processes break the spatial overlap between distinguishable and indistinguishable resident carriers, allowing the distinguishable carrier to bind to a photoexcited e-h pair. In emission, the shakeup process reunites distinguishable and indistinguishable resident carriers after the recombination.
We have self-consistently analyzed the various trends of energy shifts and broadenings of neutral excitons in their ground and excited states. The stronger energy blueshift and broadening of suboptimal excitons originate from a shakeup process that evacuates distinguishable resident carriers from the immediate vicinity of the exciton. The neutral exciton shows resilience in the absence of distinguishable resident carriers (i.e., smaller energy shift, weaker broadening, and stronger amplitude), rendering the exciton an optimal complex, and especially in moiré heterobilayer systems wherein its resonance can also become distinct. Finally, we have elucidated the intimate connection of the compressibility of the resident carrier gas to the amplitude and energy shift of the neutral exciton resonance.
Quantifying and elucidating the details of shakeup processes are interesting directions for future investigations, testing whether collective spin excitations such as magnons in valley-polarized systems or collective charge excitations such as long-wavelength plasmons can take part in these processes. Similarly, further studies are needed to quantify the compressibility of the resident carriers in moiré heterobilayers from the measured energy shift and amplitude of exciton resonances. Hopefully, the pervading nature of the Pauli exclusion principle along with the ensuing concepts of optimality of excitonic complexes, distinguishability of the e-h pair, distinctiveness of the optical resonances, and compressibility of the resident carrier gas will become the accepted approach to interpret optical studies in electrostatically-doped semiconductors.
Acknowledgements.
This work was mainly supported by the U.S. Department of Energy, Basic Energy Sciences, Division of Materials Sciences and Engineering under Award No. DE-SC0014349. Scott Crooker acknowledges support from the U.S. Department of Energy “Science of 100 T” program. The NHMFL is supported by National Science Foundation Grant No. DMR-1644779, the state of Florida, and the U.S. Department of Energy. Cedric Robert and Xavier Marie are supported by the Agence Nationale de la Recherche under the program ESR/EquipEx+ (Grant No. ANR-21-ESRE- 0025) and ANR projects ATOEMS and IXTASE.
Appendix A Hexcitons and oxcitons
Much of the way we perceive excitons in semiconductors is attributed to the seminal work of Elliott, who tailored the hydrogen model to calculate exciton states [63]. Employing the semiconductor’s dielectric constant and effective masses of its CB electron and VB hole, the two-body model successfully predicts the energy spectrum of charge-neutral excitons when using appropriate Coulomb potential models [65, 64, 66]. Similar analogy can be drawn between a trion and hydrogen ion, H-, where the trion complex is modeled as a three-body problem [67, 68, 69, 70, 71]. Solutions of this problem yield binding energies which agree with the measured energy difference between the trion and exciton resonances in the spectra of TMD monolayers [72, 73, 74, 75, 76, 77, 78].
If we wish to generalize this idea further to consider larger excitonic complexes, the following subtlety should be borne in mind. The hydrogen ion $H^{-N}$, where $N$ is the number of extra electrons, is unstable when $N>1$. One proton cannot hold more than two electrons, and therefore $H^{-N}$ dissociates to $H^{-}$ and $N-1$ electrons that end up far away from the ion and from each other. On the other hand, the stability of an excitonic complex is not a problem when dealing with an electrostatically-doped semiconductor that has $N$ Fermi seas. Here, the photoexcited pair acts as a charge-neutral hydrogen atom, and it bind to $N$ electrons with different spin-valley configurations which are near each other before excitation. To explain this point, let us consider a semiconductor whose CB has $N=2$ populated valleys and the photoexcited pair is distinguishable, as shown in Fig. 2(d). The average distance between nearby electrons, one from the valley at $K$ and the other from $-K$, is about $k_{F}^{-1}$ where $k_{F}$ is the Fermi wavenumber. When the electron density is large enough, the hexciton formation means that the photoexcited pair binds to these electrons which are already near each other at equilibrium. The same principle holds true when $N=3$ (oxciton) [24] or in multivalley semiconductors with $N\geq 4$.
Appendix B LL quantization of hexciton and oxciton resonances in a strong magnetic field
Figure 15 highlights the spectral window that pertains to the hexciton ($H$) and oxciton ($O$) resonances in Fig. 4(c). Figure 15(a) shows that the crossover from distinct hexciton to oxciton resonances at $V_{g}\gtrsim 3.5$ is corroborated by the reflectance signal with helicity $\sigma_{-}$ in Fig. 15(b), which shows that its hexciton signal jumps to the next LL at higher energy [8]. The left-pointing black arrows in Fig. 15(a) reflect the increased binding energy of the oxciton complex compared with that of the hexciton. On the other hand, the right-pointing black arrow with a yellow asterisk in Fig. 15(b) signifies the onset of Pauli blocking where the hexciton resonance becomes indistinct (the top valley of $-K$ starts to be filled).
Unlike the absorption resonances of exciton and trion complexes in Figs. 4(c) and (d), the resonances $H$ and $O$ have multiple LL branches at a given electron density, as highlighted in Fig. 15. This unique behavior was explained in Ref. [24] by considering the structure of the hexciton (or oxciton), which consists of a dark trion in its core and satellite electron(s) from the optically-active top valleys. The VB hole strongly binds to the two electrons from the bottom valleys because they are $\sim$30heavier than electrons of the top valleys [46]. The implication is that the fast relative motion of an electron in a tightly bound trion cannot be quantized in LLs like that of a free electron. On the other hand, the relative motion of a satellite electron is slower, and as such, it has larger resemblance to the LL quantized motion of a free electron.
This behavior supports the observed LL branches in Fig. 15(b) and the observation that the resonance $H_{i}$ continues to redshift despite being indistinct (Sec. IV.1). Namely, the photoexcitation does not require a shakeup process to accommodate the Pauli exclusion principle if the optically-active electron is promoted to an empty LL level and its quantized motion resembles that of a free electron.
Figure 15: Hexciton and oxciton resonances in electron-doped WSe2 when the applied magnetic field is 20 T. Taken from Ref. [22]. The bottom diagrams show corresponding absorption processes for each of the helicity-resolved optical transitions.
Appendix C Alternative explanation to the universal energy redshift
In Refs. [24, 39] , we have attributed the sustained energy redshift of hexcitons when the charge density increases to the compression of Coulomb holes that surround the hexciton. We did not pursue this explanation in the main text of this work because it raises questions when we examine Fig. 4(c). Explicitly, the energy redshift of the resonance $O$ is similar to that of the resonance $H$, and both shifts are similar in magnitude to the redshift of the resonance $X^{+}$ (but with opposite slope). Since the energy redshifts of $X^{+}$, $H$ and $O$ have similar magnitudes while their underlying excitonic complexes include one, two and three Coulomb holes, respectively, the compression of Coulomb holes is not a convincing explanation.
Appendix D Quasi-localization
There are several supporting observations that strengthen the hypothesis of quasi-localization of resident carriers at small charge densities in TMD monolayers. First, interfacial fluctuations are inevitable in every sample due to disorder and imperfections in the monolayer or its vicinity. Such fluctuations act as shallow potential wells that trap resident carriers [79, 80]. This effect fades away when (i) the charge density is large enough to screen out these traps [81], or (ii) when the cyclotron radius due to a strong magnetic field is smaller than the radius of a typical shallow trap (that is, when the cyclotron energy is larger that the binding energy to the trap) . Second, the estimated diffusion constants of thermalized singlet and triplet trions in hBN-encapsulated WSe2 monolayer are of the order of 1 cm2/s [82], whereas the diffusion constant of excitons is found to improve from 2 cm2/s to 40 cm2/s when traps are screened out [83]. Third, if triplet and singlet trions were itinerant, they would readily form hexcitons by binding to distinguishable electrons from the other bottom valley. Yet, we do not observe hexciton emission at small electron densities in WSe2 (or WS2) monolayers. Finally, quasi-localization of the resident carriers at small charge densities supports the observation that trions hardly shift in energy [62, 61], in accord with the behavior of $X^{-}_{S,T}$ in Fig. 9.
Appendix E Exchange scattering
Figure 16: Same as in Fig. 12(c) and (d), but with adding the energy shift and FWHM of the $2s$ neutral exciton resonance (dotted lines are its results at zero magnetic field). The energy shift and FWHM of each resonance are shown in a range of charge densities at which the resonance is still identifiable in the reflectance spectra.
Figure 16 shows the energy shift and FWHM of exciton resonances as a function of hole density in a hBN-encapsulated WSe2 monolayer. The results are the same as in Fig. 12, where here we add the corresponding results of the first excited state of the neutral exciton ($2s$). This resonance emerges $\sim$130 meV above that of the ground state exciton ($1s$) [55, 64]. We notice that the FWHM and energy shifts of the $2s$ excited state are stronger than those of the $1s$ ground state, which we attribute to the larger exciton size in the $2s$ state [55]. Simply put, excitons with larger radii are more sensitive to charge density changes in their vicinity [61], and to create a larger exciton mandates kicking out resident carriers from a larger region.
More interestingly, Fig. 16 shows that optimality only affects the blueshift and broadening of the exciton resonance in its ground state. In more detail, the discussion in the main text explained why the resonance of an optimal exciton shifts and broadens less in energy, evidenced by comparing the behavior of $X^{0}_{\text{1s}}(\sigma_{\pm})$. Yet, the corresponding lines that are labeled by $X^{0}_{\text{2s}}$ show no regard to whether the exciton is distinguishable (solid line; $\sigma_{-}$ at 9 T), optimal (dashed line; $\sigma_{+}$ at 9 T), or neither (dotted line; 0 T). To settle this apparent contradiction we should consider that excited state excitons are subjected to ultrafast energy relaxation through exchange scattering. As shown in Ref. [18], an exciton created in the $2s$ state is turned to a hot exciton in the $1s$ state as fast as 0.1 ps upon exchange scattering with a resident carrier when the charge density is $\sim 10^{11}$ cm-2. Such timescales reinforce the larger broadening of excited-state exciton resonances in Fig. 16. Whether the resident carrier with which the exciton interacts is distinguishable or indistinguishable determines if the exchange interaction is attractive or repulsive [45]. Yet, the energy relaxation rate is similar in both cases because the square amplitude of the exchange-scattering matrix element is indifferent to this change in sign [18]. As a result, the rate at which an exciton can relax from an excited state to the ground state depends on charge density and not on distinguishability of the resident carriers [84].